### v2

# Radio-frequency assisted switching in perpendicular magnetic tunnel junctions

Abstract
Abstract: Spin-transfer torque magnetic random-access memory (STT-MRAM) relies on nanoscale magnetic tunnel junctions (MTJs) as its fundamental building blocks. Next-generation STT-MRAM requires strategies that simultaneously improve switching energy efficiency and device endurance. Here, we present the first study of perpendicular STT-MRAM writing assisted by radio-frequency (RF) spin torque. We show that applying a small-amplitude RF pulse prior to a direct-current (DC) writing pulse enhances the MTJ switching probability, with the efficiency gain increasing at lower RF frequencies. This RF+DC writing scheme enables shorter DC pulses, thereby improving device endurance. Analytical and numerical modeling qualitatively reproduces the experimental trends, while quantitative discrepancies indicate that realistic MTJ properties beyond idealized models play an important role in RF-assisted switching.
preprint: APS/123-QED
Introduction
STT-MRAM is a leading candidate for next-generation nonvolatile memory, combining sub-nanosecond switching [37, 26, 13, 54], CMOS compatibility [35, 2, 53, 22], scalability to future technology nodes [51], and radiation hardness [31]. Significant progress has been made toward commercialization, yet the technology is currently limited to niche applications [40]. Advancing the technology to mainstream implementation requires overcoming the key challenges of relatively low energy efficiency, high write current density, and limited endurance [53]. In perpendicular MTJs (p-MTJs), which constitute the core of contemporary STT-MRAM, the write current density must be kept high enough to guarantee reliable switching, but this comes at the cost of increased power consumption and the risk of tunnel barrier breakdown [56]. Scaling the MTJ size below 20 nm further exacerbates this trade-off by making the thermal stability and write reliability harder to balance [33].
The relatively high voltage pulses applied during the write operation accelerate degradation of the ultrathin MgO barrier, leading to reduced device lifetimes. This reliability challenge stresses the need for writing schemes that deliver switching at reduced voltage levels, without compromising retention or write error rates. An alternative, spin-orbit torque MRAM (SOT-MRAM), addresses the problem of dielectric breakdown by separating the read and write current paths [30, 24, 55, 25]. Although this decoupling enables fast switching without stressing the MgO barrier, the three-terminal geometry of SOT-MRAM imposes a large footprint that constrains storage density and integration. Therefore, innovation in the engineering of the STT-MRAM stack, device design, and write protocols is crucial for the development of energy-efficient STT-MRAM.
A straightforward strategy to reduce the write voltage in STT-MRAM is to increase the duration of the DC write pulse. However, this approach relying on thermally-assisted switching comes at the cost of higher write energy and reduced energy efficiency [3, 46]. The tradeoff between the write voltage amplitude and energy efficiency motivates the search for alternative write schemes [23] that achieve reliable switching without compromising endurance or efficiency.
An attractive approach is to combine the DC write pulse with an RF current pulse [17, 11] – a method inspired by studies of microwave-assisted magnetization reversal in magnetic media [49, 52]. This RF-assisted strategy has previously been studied in in-plane MTJs and in all-metallic spin valves, where an enhanced switching probability was observed when the RF drive frequency was tuned to the ferromagnetic resonance (FMR) of the free layer (FL) [14, 18]. In thermally unstable, superparamagnetic in-plane MTJs, a strong enhancement of random telegraph switching has been demonstrated at sub-GHz frequencies, where RF excitation drives low-dimensional chaotic dynamics of the FL [32].
[…]
Figure 2: Waveforms for current-induced p-MTJ switching. a The waveform used for measurements of p-MTJ switching probability $\mathcal{P}$ in the conventional DC write pulse switching regime. b $\mathcal{P}$ as a function of DC pulse duration $t_{\mathrm{DC}}$ for three pulse amplitudes measured using the pulse sequence in a for a $D=45$ nm device. c The waveform used for measurements of RF-assisted switching.
Despite these promising results, RF-assisted switching [8, 48] in p-MTJs forming the basis of modern STT-MRAM technology has only been studied for RF frequencies near FMR [47]. In this work, we report experimental and theoretical studies of RF-assisted switching in thermally stable p-MTJs over a broad RF frequency range. We show that a small-amplitude RF pulse applied just before the DC write pulse can significantly increase the switching probability. This RF+DC writing protocol enables a reduction of the DC pulse duration, thereby improving device endurance.
Furthermore, we find that the efficiency of RF assistance increases as the RF frequency is reduced. This is a notable technological advantage, since implementing sub-GHz RF drive circuitry in CMOS is simpler and more cost-effective than realizing GHz-frequency sources required for the FMR-frequency operation. We present measurements of RF-assisted switching across multiple p-MTJ diameters and compare these experimental results to an analytical model and numerical simulations.
Results
Device characterization
Figure 3: RF-assisted increase in switching probability $\Delta\mathcal{P}$ versus RF power $P_{\mathrm{RF}}$ and RF-DC delay $\tau$. In these measurements, the RF pulse frequency $f_{\mathrm{RF}}=0.1$ GHz, and duration $t_{\mathrm{RF}}=30$ ns are fixed. The DC write pulse parameters for each p-MTJ are adjusted to achieve $\mathcal{P}=0.5$ in the absence of the RF pulse. Four p-MTJ diameters are studied: a $D=85$ nm, b $D=65$ nm, c $D=45$ nm, and d $D=25$ nm.
[…]
RF-assisted switching measurements Figure 2a illustrates the pulse sequence used for measurements of p-MTJ switching probability $\mathcal{P}$ ($0\leq\mathcal{P}\leq 1$) in the conventional DC write pulse switching regime. In this paper, we study switching from the low-resistance parallel (P) to the high-resistance antiparallel (AP) state at room temperature in zero external magnetic field. A large-amplitude square DC write pulse is applied to the device in the P state, followed by the resultant resistance readout and a longer reset pulse returning the MTJ back to the initial P state. The measurement is repeated $10^{5}$ times to reliably measure $\mathcal{P}$ for a given write pulse duration $t_{\mathrm{DC}}$ and amplitude $V_{\mathrm{DC}}$. Figure 2b shows the results of our measurements for a $D=45$ nm device as a function of $t_{\mathrm{DC}}$ for three values of $V_{\mathrm{DC}}$. All voltage values reported in this paper are voltages applied across the MTJ (Supplementary Note 1).
For studies of RF-assisted switching, the DC write pulse is preceded by an RF pulse of duration $t_{\mathrm{RF}}$, frequency $f_{\mathrm{RF}}$ and root mean square (RMS) amplitude $V_{\mathrm{RF}}$ as shown in Fig. 2c. A variable delay $\tau$ is introduced between the end of the RF pulse and the onset of the DC pulse; negative delay $\tau<0$ corresponds to temporal overlap of the RF and DC pulses. We study both the $\tau\geq 0$ and $\tau<0$ regimes of RF-assisted switching. The phase of the RF pulse in our measurements is random (Supplementary Note 1 and Figure S1).
For these measurements, we first select the values of $V_{\mathrm{DC}}$ and $t_{\mathrm{DC}}$ that give $\mathcal{P}=0.5$ in the absence of an RF pulse as determined from the type of data in Fig. 2b for each device. We then add an RF pulse with $t_{\mathrm{RF}}>t_{\mathrm{DC}}$ before the DC pulse and measure the increase in the switching probability $\Delta\mathcal{P}$ above the $\mathcal{P}=0.5$ baseline.
Figure 3 shows $\Delta\mathcal{P}$ measured as a function of RF power $P_{\mathrm{RF}}$ and delay $\tau$ for four device diameters. In these measurements, $f_{\mathrm{RF}}=0.1$ GHz, $t_{\mathrm{RF}}=30$ ns, and $P_{\mathrm{RF}}$ in dBm is defined as the power dissipated by the sample during the RF pulse:
$P_{\text{RF}}=10\cdot\log_{10}\left(\frac{V_{\text{RF}}^{2}}{R_{\text{P}}}\cdot\frac{1}{1\,\text{mW}}\right),$ (1)
where $R_{P}$ is the P state resistance (Supplementary Note 1).
Figure 4: RF-assisted increase in switching probability $\Delta\mathcal{P}$ as a function of the RF-DC delay $\tau$ at $V_{\mathrm{RF}}=0.195$ V. Four p-MTJ diameters are studied: a $D=85$ nm. b $D=65$ nm. c $D=45$ nm. d $D=25$ nm.
Figure 5: RF-assisted increase in switching probability $\Delta\mathcal{P}$ as a function of RF frequency for $\tau=-2$ ns. Error bars are described in Supplementary Note 1.
We also study the effect of time delay between the RF and DC write pulses on $\Delta\mathcal{P}$. Figure 4 shows the dependence of $\Delta\mathcal{P}$ on $\tau$ at $V_{\mathrm{RF}}=0.195$ V for four MTJ diameters. $\Delta\mathcal{P}$ monotonically increases with decreasing $\tau$ and is generally positive in both the $\tau\geq 0$ and $\tau<0$ regimes. For $\tau=-7$ ns, $\Delta\mathcal{P}$ is 0.31, 0.24, 0.27 and 0.15 for the $D=85$, 65, 45, and 25 nm devices, respectively. While at $\tau=0$, $\Delta\mathcal{P}$ is 0.13, 0.08, 0.08 and 0.06 for the $D=85$, 65, 45, and 25 nm devices, respectively.
Figure 5 shows the dependence of $\Delta\mathcal{P}$ on $f_{\mathrm{RF}}$ for the $D=45$ nm device at $P_{\mathrm{RF}}=-21$ dBm (see Supplementary Note 1 and Table S1) and $\tau=-2$ ns. These data, taken at frequencies well below the FL FMR frequency (12 GHz, Supplementary Note 2 and Figure S3), reveal that RF-assisted switching becomes more efficient with decreasing RF frequency. A qualitatively similar dependence of $\Delta\mathcal{P}$ on $f_{\mathrm{RF}}$ is observed in the $\tau\geq 0$ regime.
Theoretical results
Figure 6: Simulations of $\Delta\mathcal{P}$ as a function of normalized RF power $\tilde{P}_{\mathrm{RF}}$ and RF-DC delay $\tau$. $\Delta\mathcal{P}$ is represented by the color scale, and the insets show $\Delta\mathcal{P}$ versus $\tau$ at $\tilde{P}_{\mathrm{RF}}=-5$ dB. In these simulations, $f_{\mathrm{RF}}=0.1$ GHz and $t_{\mathrm{RF}}=30$ ns. a Macrospin simulations with the RF phase constant $\phi=0$. b Micromagnetic simulations with $\phi=0$. c Macrospin simulations with $\phi$ uniformly distributed over $\pi/2,\pi/2]$. d Macrospin simulations with $\phi$ uniformly distributed over $\pi,\pi]$.
In order to understand the mechanisms of the experimentally observed RF-assisted switching, we made room-temperature ($T=300$ K) numerical macrospin and micromagnetic simulations of the switching process. Simulations were made for all the MTJ diameters considered in the experiments (Supplementary Note 3). Here we show the results for the device with $D=45$ nm. Simulations of an ensemble of multiple replicas were performed to compute the switching probability. The magnetization dynamics of each replica were computed using both micromagnetic and macrospin solvers (see Supplementary Note 3). For the micromagnetic analysis, $10^{2}$ replicas were simulated, while for the macrospin simulations $10^{4}$ replicas were used. Both the micromagnetic and the macrospin simulations were performed using Magnetization Geometrical Integration Code (MaGICo, available at http://wpage.unina.it/mdaquino/index_file/MaGICo.html). The switching probability was then determined by counting the number of replicas that showed P to AP switching.
For these simulations, we first select the values of $V_{\mathrm{DC}}$ and $t_{\mathrm{DC}}$ that produce $\mathcal{P}=0.5$ in the absence of an RF pulse for each device. We then add an RF pulse preceding the DC pulse and calculate the increase in the switching probability $\Delta\mathcal{P}$ above the $\mathcal{P}=0.5$ baseline. Figure 6 shows the simulated $\Delta\mathcal{P}$ due to the RF pulse assist versus normalized RF power $\tilde{P}_{\mathrm{RF}}$ and delay $\tau$ between the RF and DC write pulses. The RF power is normalized to the minimal theoretical power required to switch the magnetic state with a DC current in the absence of thermal fluctuations ($T=0$ K) [41]. This minimal power is analytically calculated in the macrospin approximation for the values of the FL dimensions, the uniaxial magnetocrystalline anisotropy constant $K_{\mathrm{AN}}$, saturation magnetization $M_{\mathrm{s}}$, and Gilbert damping constant $\alpha$ given in Table I of Methods.
Figure 7: Analytic model predictions of $\Delta\mathcal{P}$ as a function of normalized RF power $\tilde{P}_{\mathrm{RF}}$ and RF-DC delay $\tau$. a $\phi=0$. b $\phi$ uniformly distributed over $\pi/2,\pi/2]$. c $\phi$ uniformly distributed over $\pi,\pi]$.
Figure 6a shows macrospin $\Delta\mathcal{P}$ as a function of $\tilde{P}_{\mathrm{RF}}$ and $\tau$ when the phase constant $\phi$ of the RF voltage $\sqrt{2}V_{\mathrm{RF}}\cos(\omega_{\mathrm{RF}}t+\phi)$ is set to zero. Notice that $t=0$ corresponds to the time instant where the RF pulse is applied and $\omega_{\mathrm{RF}}=2\pi f_{\mathrm{RF}}$. These results are significantly different from the experiment in Fig. 3c and Fig. 4c. This is most clearly seen from the comparison of the inset in Fig. 6a showing $\Delta\mathcal{P}$ versus $\tau$ at $\tilde{P}_{\mathrm{RF}}=-5$ dB and the experimental data in Fig. 4c. Indeed, the measured $\Delta\mathcal{P}$ is a monotonically decreasing function of $\tau$, while the simulated $\Delta\mathcal{P}(\tau)$ has a peak at $\tau=-3$ ns. The micromagnetic results in Fig. 6b are similar to the macrospin with a slightly enhanced $\Delta\mathcal{P}$. The number of replicas in the macrospin simulations is large enough to yield a $\sim$1 % error in $\Delta\mathcal{P}$ while the micromagnetic results are precise within $\sim$10 % (Supplementary Note 3).
The discrepancy with the experiment is expected because the simulations assume a fixed phase $\phi=0$, while the measurements average over random phases. To address this discrepancy, we performed macrospin simulations with $\phi$ uniformly distributed over $\pi/2,\pi/2]$ (Fig. 6c) and $\pi,\pi]$ (Fig. 6d) intervals, the latter directly mimicking the experimental conditions. The inset in Fig. 6d is qualitatively similar to the corresponding experimental results in Fig. 4c, although quantitative discrepancies remain. Similar results were obtained in micromagnetic simulations and for devices with other studied MTJ diameters (Supplementary Note 3 and Figure S4).
The numerically calculated dependence of $\Delta\mathcal{P}$ on $\tilde{P}_{\mathrm{RF}}$ and $\tau$ can be further elucidated by analytical modeling in the framework of damping-switching dynamics [27, 15]. This analytical model makes five assumptions:
(i) macrospin approximation.
(ii) prior to the RF+DC waveform, the distribution of replicas is assumed to be described by the Néel-Brown model [7] in the high-energy-barrier case.
(iii) magnetization dynamics is governed by the deterministic Landau-Lifshitz-Gilbert (LLG) equation.
(iv) for $\tau>0$, the RF pulse does not affect the switching probability.
(v) for $\tau<0$, the RF+DC waveform is replaced by an effective DC pulse of duration $t_{\mathrm{DC}}$ and amplitude $\overline{V}_{\mathrm{ST}}$:
$\overline{V}_{\mathrm{ST}}=V_{\mathrm{DC}}+\sqrt{2}V_{\mathrm{RF}}\,\frac{|\tau|}{t_{\mathrm{DC}}}\mathrm{sinc}\left(\frac{\omega_{\mathrm{RF}}\tau}{2}\right)\cos\left(\omega_{\mathrm{RF}}\left(t_{\mathrm{RF}}+\frac{\tau}{2}\right)+\phi\right)\,$, (2)
which is $V_{\mathrm{DC}}$ plus the average RF voltage over the duration of the DC pulse, with $t_{\mathrm{RF}}$ the duration of the RF pulse. The theoretical justification of these five assumptions is discussed in Supplementary Note 3. Then, the magnetization switching dynamics is a damping-switching process in a uniaxial, bistable magnet between the stable states $m_{z0}$ and $m_{z1}$, where $m_{z0}$ ($m_{z1}$) belongs to the region of stability of $m_{z}=1$ ($m_{z}=-1$). The magnetization trajectory from $m_{z0}$ to $m_{z1}$ under the action of $\overline{V}_{\mathrm{ST}}$ can be expressed in closed form [15], which can be used to map the ensemble of initial conditions to the ensemble of final states. Therefore, the stochastic nature of the magnetization process is included only in the statistical distribution of the initial conditions. The FL switches if $m_{z}$ changes sign compared to $m_{z0}$ after the write pulse. Starting with the initial distribution with $m_{z0}>0$, the switching probability can be computed by integrating the PDF over the range $m_{z}\in[-1,0]$ that defines the well of the switched magnetic state $m_{z}=-1$. This PDF solution is described in Supplementary Note 3. The switching probability $\mathcal{P}$ for $t^{\prime}=t-(t_{\mathrm{RF}}+\tau)>0$ is given by:
[…]
The quantity $\overline{H}$ is the effective DC voltage pulse amplitude normalized to $V_{\mathrm{c0}}$, which is the critical DC voltage for switching at $T=0$. The quantity $\mu_{z}^{-1}(0,t^{\prime},\overline{H})$ is the initial condition $m_{z0}$ that by the deterministic dynamics under the action of the normalized voltage $\overline{H}$ for a time period of duration $t^{\prime}$ leads to the final state $m_{z1}=0$. The quantity $k_{\mathrm{eff}}/\tilde{T}$ is the ratio between the normalized effective anisotropy and the normalized temperature. They are given by: $k_{\mathrm{eff}}=2K_{\mathrm{AN}}/(\mu_{0}M_{s}^{2})-D_{\mathrm{z}}+D_{\perp}$, and $\tilde{T}=k_{\mathrm{B}}T/(\mu_{0}M_{s}^{2}V)$. Here, $K_{\mathrm{AN}}$ is the uniaxial magnetocrystalline anisotropy, $D_{\mathrm{z}}$ and $D_{\mathrm{\perp}}$ are the demagnetizing factors along the symmetry axis and the perpendicular direction $(2D_{\perp}+D_{z}=1)$, $k_{\mathrm{B}}$ is the Boltzmann constant, $\mu_{0}$ is the vacuum magnetic permeability, $M_{s}$ is the saturation magnetization, and $V$ is the volume of the free layer. All the quantities are expressed in SI units. Taking into account that PDF depends on $\phi$ via $\overline{H}$, the switching probability for $\phi$ uniformly distributed over the experimentally relevant interval $\pi,\pi]$ is:
$\mathcal{P}(t^{\prime})=\frac{\int_{0}^{2\pi}\mathrm{erfi}\left(\sqrt{\frac{k_{\mathrm{eff}}}{2\tilde{T}}}\mu_{z}^{-1}(0,t^{\prime},\overline{H})\right)d\phi}{2\pi\,\mathrm{erfi}\left(\sqrt{\frac{k_{\mathrm{eff}}}{2\tilde{T}}}\right)}\,$. (4)
In Fig. 7, we use Eq. 3 and Eq. 4 to compute $\Delta\mathcal{P}$ for the same parameters as numerical results in Fig. 6. In the case of $\phi=0$ and $\phi\in[-\pi/2,\pi/2]$ (Fig. 7a,b), the agreement with the numerical results is quantitative, but its quality becomes worse for large values of $P_{\mathrm{RF}}$ and $\tau<0$, where the analytical results underestimate $\Delta\mathcal{P}$. Since the effective write pulse amplitude $\overline{V}_{\mathrm{ST}}$ is an oscillatory function of $\tau$, $P_{\mathrm{RF}}$ oscillates between positive and negative values for $\tau<0$.
In the case of the fully random phase $\phi$ distributed in $\phi\in[-\pi,\pi]$, analytical calculations in Fig. 7c do not predict an increase of $\mathcal{P}$, while numerical simulations in Fig. 6d show $\Delta\mathcal{P}>0$ for large values of $P_{\mathrm{RF}}$ and $\tau<0$. This is because $\mathcal{P}(\phi)$ in Eq. 3 is a periodic function of $\phi$, and the enhancement of $\mathcal{P}$ achieved for a given $\phi^{*}$ is approximately canceled by the reduction of $\mathcal{P}$ for the phase $\phi^{*}+\pi$. Although this cancellation is not exact, $\Delta\mathcal{P}$ resulting from the averaging in $\phi$ is small $\sim 0.01$.
In numerical simulations in Fig. 6d with sufficiently high $\tilde{P}_{\mathrm{RF}}$ and $\tau<0$, $\Delta\mathcal{P}$ is not negligibly small because a larger number of switching events with respect to those predicted by the analytical model in Fig. 7c occur in the time interval where the RF pulse adds constructively to the DC pulse voltage. This fact is not surprising since the assumption (v) of the analytical model does not hold anymore (Supplementary Note 3). As a result, numerical simulations produce higher $\Delta\mathcal{P}$ than the analytic model.
Discussion
Figure 3 reveals that RF-assisted switching enhancement is observed for both positive and negative delays $\tau$ between the RF and DC pulses. The enhancement for $\tau<0$ can be expected because the overlap between the RF and DC pulses can result in the net applied voltage exceeding $V_{\mathrm{DC}}$ in an interval of the combined RF+DC waveform. However, the RF-enhanced switching in the $\tau\geq 0$ regime is surprising. This latter regime is also important from a practical point of view because the combined RF+DC waveform voltage never exceeds $V_{\mathrm{DC}}$, leading to enhanced switching without extra stress on the junction. From another angle, introduction of an RF pulse allows one to reduce the DC pulse duration $t_{\mathrm{DC}}$ for a targeted switching probability compared to the DC-only protocol. Since longer DC write pulses are known to stress the MTJ barrier oxide and reduce device endurance [1, 19, 50], the RF+DC switching protocol in the $\tau\geq 0$ regime studied here is expected to enhance STT-MRAM endurance. Prior studies of STT-MRAM reported a nearly order-of-magnitude improvement in the MgO write cycles before the dielectric breakdown when $t_{\mathrm{DC}}$ was reduced from $10$ ns to $5$ ns [50].
The maximum RF pulse amplitude used in this study, $V_{\mathrm{RF}}=0.195$ V, is much smaller than the DC write pulse amplitude $\sim 1$ V. Therefore, in the $\tau\geq 0$ regime, this small-amplitude RF assist does not diminish device endurance and only marginally increases the write energy while increasing the switching probability. This makes RF assist at $\tau=0$ a viable strategy for an improved STT-MRAM write protocol. For example, addition of a small-amplitude RF pulse at $\tau=0$ in Fig. 4a, leads to a modest 17 % increase of the write energy while giving a 13 % increase in the switching probability without affecting the device endurance. Our calculations of the energy per write (Supplementary Note 1 and Figure S2) identify regimes in which the RF+DC and DC protocols have equal write energies, while $t_{\mathrm{DC}}$ is shorter in the RF+DC protocol. Figure S2 shows an example where the addition of a small-amplitude RF pulse reduces $t_{\mathrm{DC}}$ while maintaining the same total write energy for $\tau=-7$ ns. A comprehensive follow-up study of STT switching as a function of RF and DC pulse parameters is required to determine whether the RF+DC protocol can achieve higher energy efficiency than the DC-only write protocol. The present study focuses on regimes $\tau\geq 0$ where the RF+DC protocol improves the switching probability without compromising STT-MRAM endurance.
Figure 3 also demonstrates that the small-amplitude RF pulse increases the switching probability for all device sizes down to $D=25$ nm. This is promising for scalability of the RF+DC write. While the observed enhancement of switching probability $\Delta\mathcal{P}$ at $V_{\mathrm{RF}}=0.195$ V is smaller for the 25 nm device compared to the 85 nm device, the RF power dissipated by the 25 nm device at this RF assist voltage is 10 dB lower due to the higher resistance of the smaller MTJ. The effectiveness of the RF assist for all device sizes indicates that micromagnetic effects expected to be more pronounced in the larger devices do not play a crucial role in the observed effect. This is corroborated by our theoretical results in Fig. 6 revealing similar impact of the RF assist in the macrospin and micromagnetic approximations.
In principle, ohmic heating induced by the RF pulse can contribute to the enhancement of switching probability $\Delta\mathcal{P}$. However, because time-averaged ohmic heating is frequency independent, MTJ heating averaged over the RF pulse duration cannot account for the strong frequency dependence seen in Fig. 5. Circuit calibrations used in our experiments ensured uniform RF pulse power across all frequencies (See Supplementary Note 1 and Table S1). If time-averaged MTJ temperature increase induced by the RF pulse were the dominant mechanism, the observed switching enhancement would be comparable across all frequencies at constant power. The strong frequency dependence of the RF assist demonstrated in Fig. 5 therefore eliminates time-averaged ohmic heating of the MTJ by the RF pulse as the primary mechanism of the observed RF-induced switching enhancement [34].
Our theoretical predictions of RF-assisted switching in Fig. 6d are qualitatively similar to the experimental results in Fig. 3c and Fig. 4c. The theory captures the threshold character of the effect in the RF power and the observed decrease in the switching probability with the delay time $\tau$ between the RF and DC pulses. However, there are two quantitative differences. First, the theoretical threshold power is higher than the experimental one. Second, theoretical results shown in the insets of Fig. 6 suggest that the effect vanishes for delay times $\tau$ exceeding $\sim 1$ ns, while the observed effect in Fig. 4 persists up to $\tau\approx 6$ ns.
The theoretically predicted $\sim 1$ ns time scale is reasonable because it is similar to the characteristic magnetic energy relaxation time of the FL $\tau_{E}=1/\left(4\pi\alpha f_{\mathrm{FMR}}\right)=0.66$ ns, where $\alpha=0.0097$ and $f_{\mathrm{FMR}}=12$ GHz as determined by spin torque FMR measurements (Supplementary Note 2 and Figure S3). For $\tau\gg\tau_{E}$, all magnetic energy added to the FL magnetization by the RF pulse is expected to dissipate away, and thus magnetization dynamics driven by the DC pulse is expected to be unaffected by the RF pulse. The observation of RF assist effects for $\tau\gg\tau_{E}$ suggests that, in addition to FMR dynamics, the FL also supports significantly slower magnetic dynamics.
Signatures of slow magnetic dynamics have been previously observed in a different experiment [39] and were attributed to regions at the FL/MgO interface with reduced magnetization and weak exchange coupling. Our data in Fig. 4 provide additional support for this hypothesis of slow dynamics in the FL of a p-MTJ. Moreover, the greater enhancement of switching probability $\Delta\mathcal{P}$ at lower RF frequencies as shown in Fig. 5 suggests these dynamics are more efficiently excited at lower frequencies. This effect may be rooted in the structural, chemical, and magnetic inhomogeneity of the ultrathin CoFeB FL found in MTJ devices [5]. Electrical transport, FMR and magnetometry studies on ultrathin films and MTJ nanopillars based on Ta/CoFeB/MgO multilayers suggest that iron-based magnetic oxides form at the CoFeB/MgO interface [5, 4]. These slow magnetic dynamics attributed to the FL inhomogeneity could contribute to the enhancement of switching probability $\Delta\mathcal{P}$ for $\tau\geq 0$ but are beyond the scope of the theoretical framework of this paper.
The threshold character of the RF-induced enhancement of switching in Fig. 3 suggests the importance of nonlinear magneto-dynamics [10] for the RF assist. Both the threshold character of the RF assist in the RF power and the frequency dependence of its efficiency in Fig. 5 are consistent with the prediction of magnetization reversal driven by low-dimensional chaos induced by the RF drive [32]. However, this mechanism is expected to be inactive in perfectly axially symmetric cylindrical structures with uniaxial magnetic anisotropy. This is because the dynamic system is effectively two-dimensional with one dimension represented by the magnetization polar angle and the other by the RF drive phase. According to the Poincaré – Bendixson theorem [12], such a system does not exhibit chaotic dynamics.
However, real nominally circular MTJ devices are known to deviate from this ideal picture and exhibit a weaker random in-plane uniaxial anisotropy in addition to the dominant out-of-plane uniaxial anisotropy [21, 28]. This anisotropy can arise from both the MTJ shape imperfections and from crystallographic anisotropy of the randomly oriented FL grains. In this case, both the polar and azimuthal angles of the magnetization enter the system’s Lagrange function, and the RF drive makes the system effectively three-dimensional [42, 43, 16]. In such a system, RF-driven chaotic dynamics is expected above a threshold amplitude of the drive, and the threshold magnitude decreases with deceasing frequency. This RF-driven chaos mechanism can contribute to the observed RF assist and reduce the threshold for the onset of RF-induced switching over that theoretically predicted for ideal cylindrically-symmetric devices in Fig. 6.
A potential mechanism of switching probability enhancement not included in the present theoretical description is the dynamics of the FL temperature and its coupling with the magnetization switching dynamics. An RF current applied to the MTJ produces a temperature increase that has a time-independent component that also does not depend on $f_{\mathrm{RF}}$ and an oscillatory component at the frequency $2f_{\mathrm{RF}}$ with a frequency-dependent amplitude approximately given by $T_{0}/\sqrt{1+(4\pi\tau_{\mathrm{\theta}}f_{\mathrm{RF}})^{2}}$, where $\tau_{\mathrm{\theta}}$ is the thermal time constant of the system and $T_{0}$ is the temperature oscillation amplitude in the low frequency limit [44, 36]. If we assume that the frequency dependence of the RF assisted switching in Fig. 5 is linked to these time-dependent temperature oscillations, we can fit the data in Fig. 5 to $\Delta\mathcal{P}_{0}/\sqrt{1+(4\pi\tau_{\mathrm{\theta}}f_{\mathrm{RF}})^{2}}$ in order to extract $\tau_{\mathrm{\theta}}$, where $\Delta\mathcal{P}_{0}$ is the switching probability enhancement in the low frequency limit. Such a fitting procedure gives $\tau_{\mathrm{\theta}}=0.18\pm 0.01$ ns.
Such a small thermal time constant is an order of magnitude below previously reported values of $\tau_{\mathrm{\theta}}$ for nanoscale MTJs [9]. In addition, $\tau\gg\tau_{\theta}$ is inconsistent with the data in Fig. 4, which shows that the effects of the RF pulse on the switching dynamics persists for nanoseconds after the end of the RF pulse. Note that the RF pulses of the magnitude and duration used in this study do not produce any measurable MTJ switching on their own without a follow-up DC pulse.
We, therefore, conclude that while a contribution of Ohmic heating to the observed RF assist cannot be completely excluded, it cannot be a simple process governed by a single time constant. Investigating a more complex mechanism would require careful measurements of the thermal time constants together with the development of a theory describing coupled thermal and magnetization dynamics. This analysis is beyond the scope of the present work and will be addressed in future studies.
Methods
Sample description The bottom electrode/ SAF/ MgO/ CoFeB(2.21 nm)/ top electrode multilayer stack was deposited by magnetron sputtering and patterned into circular MTJs using e-beam lithography and ion mill etching. Saturation magnetization of the CoFeB FL was determined by vibrating sample magnetometry, while its out-of-plane uniaxial magnetic anisotropy and Gilbert damping were determined by ST-FMR measurements (Supplementary Note 2) as summarized in Table 1.
Table 1: Material parameters of the MTJ CoFeB free layer: thickness, saturation magnetization $M_{\mathrm{s}}$, uniaxial magnetocrystalline anisotropy constant $K_{AN}$, Gilbert damping constant $\alpha$.
Quantity Value Unit
FL thickness $2.21$ nm
$M_{s}$ $1039$ $\mathrm{kA/m}$
$K_{\mathrm{AN}}$ $8.6\times 10^{5}$ $\mathrm{J/m^{3}}$
$\alpha$ $0.0097$
Spin torque ferromagnetic resonance We use ST-FMR to measure the frequency and linewidth of the FL spin wave modes as a function of out-of-plane magnetic field, and extract the FL Gilbert damping constant $\alpha$ and the uniaxial magnetocrystalline anisotropy constant $K_{AN}$ at room temperature from these data [20, 38, 6]. The values presented in Table 1 are an ensemble average of multiple devices. In these measurements, a microwave voltage is applied to an MTJ via the RF port of a bias tee. The applied voltage induces an RF spin transfer torque that excites spin waves in the FL and leads to oscillations of the sample resistance at the drive frequency due to tunneling magneto-resistance of the MTJ. These resistance oscillations mix with the RF current through the MTJ and generate a rectified DC voltage, which is measured by a lock-in amplifier through the DC port of the bias tee. The magnetic field modulation technique is employed to improve the signal-to-noise ratio [20].
RF-assisted switching probability measurements Current-induced switching of the MTJ FL is observed by monitoring the change in the MTJ resistance in a circuit where the MTJ forms a voltage divider with a fixed resistor. We bias the voltage divider by a low-level DC voltage and use a 16-bit 2 MHz data acquisition board (DAQ) to measure time-dependent voltage across the MTJ. To switch the FL, we apply a DC voltage pulse preceded by an RF voltage pulse to the MTJ. Details of the circuit and the waveform protocol are given in Supplementary Note 1. Following the RF+DC write waveform, we use DAQ to measure the resultant resistance state of the MTJ, which registers either switching or non-switching. A lower-amplitude, longer-duration DC reset pulse of the opposite polarity to the DC write pulse is then used to return the device to its initial state, and the DAQ is used to confirm the successful reset. The RF pulse is generated by using a broadband microwave mixer connected to a microwave generator controlling the RF frequency and a square pulse generator controlling the output RF pulse duration. The amplitude of the RF pulse is controlled by both generators. All three pulse generators used in the measurements are synchronized with each other and the DAQ as described in Supplementary Note 1. The microwave generator operates in continuous wave mode, resulting in a random phase $\phi$ of the RF pulse. The delay $\tau$ between the RF and DC pulses is controlled through a combination of the cable length and a programmable internal delay of the DC write pulse generator. The switching probability $\mathcal{P}$ is determined as a ratio of the number of successful switching events to the total number of switching attempts. The voltage levels of all RF+DC waveforms used in the measurements are characterized using a high-bandwidth real-time oscilloscope.
Data availability
All data generated or analyzed during this study are included in this published article and are available from the corresponding author upon reasonable request.
Acknowledgements
I.N.K. and P.M.B. thank Jordan Katine for stimulating discussions.
This work was primarily supported by Western Digital Technologies (grant number: not applicable). Funding from the National Science Foundation via awards ECCS-2213690 and DMREF-2324203 as well as from the University of California National Laboratory Fees Research Program (grant number: not applicable) is also acknowledged. The authors thank the Eddleman Quantum Institute (EQI) for partial support of this work (grant number: not applicable). S.P gratefully acknowledges financial support for this publication by the Fulbright Visiting Scholar Program, which is sponsored by the U.S. Department of State and the Italian Ministry of Foreign Affairs and administered by the U.S. - Italy Fulbright Commission (grant number: not applicable). The contents of this publication are solely the responsibility of the author and do not necessarily represent the official views of the Fulbright Program, the Government of the United States, or the U.S. - Italy Fulbright Commission.
Author contributions
P.M.B., W.J. and C.D. made the samples. M.H. and I.N.K. designed the experiment. M.H. made the measurements. S.P., M.d’A. and C.S. developed the theoretical model. S.P. made numerical simulations. P.M.B and I.N.K. conceived the idea and supervised the project.
Competing interests
The authors declare no competing interests.